# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ 4.1 Models
We evaluated several publicly-available models on Dynamic-SUPERB Phase-2: SALMONN, LTU-AS, Qwen-Audio, Qwen2-Audio, WavLLM, MU-LLaMA (Liu et al., 2024), GAMA (Deshmukh et al., 2024). SALMONN is further categorized into 7B and 13B versions based on the size of its LLM component. All of these models are publicly available, and we utilized their official implementations for inference without any modifications.
The first five models are instruction-based speech models, each also trained with audio understanding capabilities. However, as they were not trained on music data, we do not expect them to perform well on music-related tasks. Hence, we further included MU-LLaMA and GAMA. MU-LLaMA is specifically designed for various music-understanding tasks, and GAMA is trained with both general audio data and a music corpus. Similar to the speech models, MU-LLaMA and GAMA adopt an LLM-based framework. They use a music or audio encoder, such as MERT (LI et al., 2024) or Q-former, to convert music or audio into features, which the LLM then uses as prompts for subsequent reasoning. The sampling strategies used for generating outputs from the LLMs were retained as defined in their official implementations. Finally, we implemented a cascaded system baseline called Whisper-LLaMA. The system first transcribes audio using Whisper-v3-large, and then LLaMA3.1-8B processes the transcriptions to perform various tasks based on the provided instructions.
When evaluating models on these diverse tasks, several challenges inevitably arise. One challenge was testing the baseline models on tasks that involve multiple audio inputs. For example, in speaker verification, a model determines whether two utterances are produced by the same speaker. Among the evaluated models, only Qwen-Audio and Qwen2-Audio provide interfaces that allow inputting multiple audio files. For the remaining models, we concatenated all audio files, separated by 0.5 seconds of silence, and used the concatenated file as input.
Last, different models have their own maximum supported audio durations. Models using Whisper to encode speech face an inherent limitation: Whisper can process audio files of up to 30 seconds, and some models are restricted to handling even shorter durations. Consequently, we retained the original model settings and did not modify the model architecture to accommodate longer audio clips. A preliminary analysis of all audio (Appendix G) in Dynamic-SUPERB Phase-2 shows that only a small proportion (around 7%) exceeds 30 seconds in length. Hence, we believe our settings do not largely impact the evaluation results and ensure reasonable inference efficiency.
@@ 4.2 Evaluation metrics
The outputs of universal speech models are natural language sentences, which makes it difficult to directly assess their correctness using conventional metrics. For classification tasks, such as emotion classification, a conventional task-specific model outputs a label from predefined emotions in the dataset, while a universal model generates sentences like “The speaker sounds happy.” In this case, we can easily assess the correctness of the former by comparing labels, but this approach is not applicable to the latter. Similarly, in regression tasks, a natural language response like “Using MOS scoring criteria, I give this audio a score of 3 out of 5” makes it difficult to directly use previous evaluation metrics such as mean square error.
@@
In evaluation, we categorize tasks into three types: (1) classification, (2) regression, and (3) sequence generation. For each type, we use different pipelines to evaluate the models’ outputs. For classification tasks, we employ external LLMs (GPT-4o) as referees, with the temperature set to be zero for evaluation consistency, to evaluate whether the outputs from speech or music models222To avoid ambiguity, unless specified otherwise, the terms ‘output” or model output” in this section refer to the results generated by the models in Sec. 4.1. match the ground truth. This approach has been widely adopted in the NLP community (Wang et al., 2023; Liu et al., 2023b; Chiang & Lee, 2023), and we extend its application to speech research. We design a prompt that includes the task instructions, the model’s output to be evaluated, and the corresponding ground truth. The LLM judge gets this prompt and determines if the output aligns with the ground truth using a chain-of-thought reasoning process. By leveraging the strong instruction-following capabilities of these LLMs, we constrain them to provide the final decision in a fixed format, allowing us to extract the answer using simple methods such as regular expressions333We have tried to ask the speech universal models to directly output their decisions in a fixed format, but they do not always follow the instructions. Therefore, the evaluation pipeline described here is necessary.. We then define accuracy as the percentage of outputs that the LLM judge considers aligned with the ground truth. Please refer to Appendix E for details on the prompts and the alignment between LLMs and humans.
@@
For regression tasks, the aforementioned method cannot be directly applied because performance is not assessed using hard labels. Instead, we employ LLMs (GPT-4o) as a post-processor to transform natural language responses into a format compatible with the original metrics used in these tasks. Specifically, we developed an additional prompt that directs the LLMs to extract essential information from the output and convert it into the same format as the ground truth. If the output lacks any correct or relevant information that can be converted to match the ground truth format, the LLM returns “N/A” to mark it as invalid. Since conventional metrics cannot accommodate “N/A” and setting a default value is unreasonable due to the variability of different metrics, we also calculate the N/A rate for each regression task, defined as the percentage of invalid outputs within a task. A higher N/A rate indicates that the model struggles with following instructions, which may indirectly affect its performance on the task.
@@ 5.1 Domain-Level Performance Comparison in Taxonomy
Figure 3 presents a domain-level comparison of relative scores across different models. Whisper-LLaMA consistently has zero scores throughout all domains, serving as the reference baseline for evaluating other models. Starting with speech domains, no model outperforms Whisper-LLaMA in speech recognition and spoken language understanding. This highlights that using ASR such as Whisper remains a strong baseline for language understanding, as text more explicitly represents semantic information than speech.
Conversely, in the speaker and paralinguistics domains, all models surpass the baseline. This improvement occurs because the ASR process in the cascaded system tends to discard critical information from the speech signal, such as speaker characteristics, pitch, and emotion. In contrast, other models utilize soft representations that retain more speech information, giving them the potential to learn beyond ASR, and some of these models were also trained to perform specific tasks within certain domains 444Although the models have been trained on some tasks in the same domains, the training tasks of these models generally do not use the same instructions as Dynamic-SUPERB and do not offer as broad coverage.. Additionally, music-specific large language models like GAMA and Mu-LLaMA perform poorly on speech recognition and understanding tasks, which is expected since they were primarily designed for music understanding. Nevertheless, they surprisingly achieved scores comparable to the speech models in speaker and paralinguistics.
Turning to audio and music domains, where ASR models cannot transcribe non-speech information into text, most models outperform the baseline in several areas, such as music classification and sound event detection. Surprisingly, spoken language models primarily trained on speech data (Qwen, SALMONN, WavLLM)555Qwen and SALMONN are trained on a mixture of speech, music, and audio data, with speech data being the dominant component. outperform the two music models (GAMA and Mu-LLaMA) in various music-related domains such as harmony and pitch, music classification, and rhythm analysis. We speculate that training on diverse data, even with significant differences in signal-level characteristics, enhances performance across domains, emphasizing the importance of developing unified models for speech, music, and general audio processing.
@@ 5.2 Core tasks results
| Whisper-LLaMA | 100.1 | 36.5 | 55.5 | 28.7 | 34.0 | 46.5 | 53.6 | 51.3 | 53.0 | 1068.7 |
| SALMONN-7B | 25.4 | 30.5 | 25.5 | 39.8 | 15.1 | 49.0 | 61.6 | 61.5 | 65.5 | - |
| SALMONN-13B | 24.6 | 2.0 | 9.0 | 30.9 | 2.8 | 51.5 | 29.6 | 106.3 | 93.5 | - |
| Qwen-Audio | 100.6 | 60.5 | 15.5 | 33.7 | 69.4 | 48.0 | 42.2 | 90.1 | 47.5 | 8852.2 |
| Qwen2-Audio | 101.0 | 47.0 | 11.5 | 68.1 | 36.7 | 53.5 | 42.4 | 55.7 | 55.5 | 4664.4 |
| WavLLM | 100.0 | 43.0 | 4.5 | 79.1 | 6.9 | 49.5 | 50.8 | 84.9 | 66.5 | 3621.2 |
| LTU-AS | 102.8 | 1.0 | 0.0 | 37.2 | 96.3 | 45.0 | 10.2 | 559.3 | 40.0 | 14923.3 |
| GAMA | 100.4 | 2.0 | 0.0 | 17.7 | 116.7 | 2.5 | 2.7 | 4750.0 | 33.5 | 187.6 |
| Mu-LLaMA | 110.3 | 4.0 | 0.0 | 25.0 | 103.7 | 51.0 | 1.0 | 15869.6 | 49.0 | - |
@@
We tested the performance of the models on core tasks that are widely studied and commonly used. Table 2 presents the detailed results of the models on SUPERB tasks (please refer to Appendix D for results on HEAR and MARBLE), displaying each task’s performance in its own metric rather than relative scores. Due to space limitations, please refer to Appendix D for results on HEAR and MARBLE. It is worth noting that these tasks have been adapted to the Dynamic-SUPERB framework. Therefore, the results are not directly comparable to previous studies that tested on the original SUPERB. However, they indicate current performance levels on these tasks.
No single model excels across all tasks. In phoneme recognition (PR), the SALMONN models were the only ones to achieve a relatively lower phoneme error rate (PER) compared to the others. For keyword spotting (KS), Qwen-Audio was the only model to perform slightly better than a random guess (50%). Whisper-LLaMA surpassed all other models in intent classification (IC). Regarding emotion recognition (ER), Qwen2-Audio and WavLLM stood out, with the former even achieving almost 80% accuracy. In ASR, SALMONN-13B and WavLLM are the only two models that achieved a word error rate (WER) lower than 10%. Query-by-example (QbE) appears to be very challenging for all models, as none performed better than a random guess (50%). Slot-filling (SF) is evaluated using two metrics: F1 for slot type and CER for slot value. While Whisper-LLaMA and SALMONN-7B are the top two in this task, their results are still far from ideal. In speaker verification, SALMONN-13B was the only model to significantly outperform the others, achieving over 90% accuracy. In speaker diarization, all models performed poorly, with some even reaching a 100% N/A rate. Comparing Figure 3 and Table 2, we observed that performance on core tasks sometimes deviates from trends observed at the domain level. For example, although WavLLM excels in ER, it performs poorly in paralinguistics, revealing limited capabilities and highlighting the need to enhance its generalizability. In summary, among tested models, some excel in specific tasks, while others perform poorly across the board, and none dominate across all tasks. We believe these findings provide valuable insights for researchers aiming to improve models, starting from core tasks and progressively tackling each domain in the benchmark for broader applicability.
@@ 6 Conclusions
This paper presents Dynamic-SUPERB Phase-2, the largest benchmark for evaluating instruction-based universal spoken language models, building upon collaborative efforts across the research community. Dynamic-SUPERB covers a wide range of diverse tasks and offers a fine-grained task taxonomy. The recent models show good performance on specific tasks but poor generalization across common tasks like those in SUPERB, highlighting the need for further research on universal models. All materials will be made openly available to facilitate reproduction and benchmarking. We sincerely invite researchers to join our vibrant community and collaborate to advance the field.
@@ 7 Ethics Statement
Dynamic-SUPERB Phase-2 includes several datasets, and we asked contributors to describe how they utilized these datasets. Most contributors derived task data by uniformly sampling from the original datasets without applying any special processing. As a result, Dynamic-SUPERB Phase-2 inevitably inherits biases present in these datasets across the tasks.
@@ 8 Reproducibility Statement
As mentioned in the main text, we will open-source all task data and the evaluation pipeline to ensure reproducibility. Additionally, we set the GPT-4o temperature to 0 to stabilize the evaluation process. In Appendix E, we also present a preliminary study on using open-source LLMs for evaluation, showing promising correlations with human annotators.
@@ Appendix B Dynamic-SUPERB tasks
††footnotetext: We are still reviewing the licenses of all tasks and have identified some potential issues, which we are working to resolve. Therefore, the list and results are subject to change.| Domain | Task | Dataset |
| Audio / Singing Analysis / Vocal Techniques | Vocal Technique Classification | VocalSet (Wilkins et al., 2018) |
@@
| Audio / Singing Analysis / Lyrics | Lyric Translation | SingSet (recorded by the task contributor) |
| Audio / Singing Analysis / Lyrics | Song Lyric Recognition | SingSet (recorded by the task contributor) |
| Audio / Spatial Audio Analysis | Sound Position Prediction | BinauralSoundPerception (Vasudevan et al., 2020) |
| Audio / Spatial Audio Analysis | Audio Spatial Distance Prediction | Spatial LibriSpeech (Sarabia et al., 2023) |
| Audio / Spatial Audio Analysis | HowFarAreYou | 3DSpeaker (Zheng et al., 2023) |
| Audio / Safety / Audio Integrity | Scene Fake Detection | SceneFake (Yi et al., 2024) |
| Audio / Safety / Audio Integrity | Audio Editing Identification | Peoples Speech (Galvez et al., 2021) |
| Audio / Harmony & Pitch/Harmony | MARBLE Key Detection | Giantsteps Key (Knees et al., 2015) |
| Audio / Harmony & Pitch/Harmony | Chord Classification | Acoustic Guitar and Piano (DeepContractor, 2024) |
| Audio / Safety / Deepfake | Singing Voice Deepfake Detection | CtrSVDD ACEKiSing M4Singer (Zang et al., 2024) |
| Audio / Safety / Deepfake | Audio Deep Fake Detection | LJSpeech WaveFake MUSDB18HQ (Frank & Schönherr, ) |
| Audio / Sound Event / Advanced Understanding | Multichannel Sound Event Understanding | STARSS23 (Shimada et al., 2024) |
@@
| Audio / Sound Event / Advanced Understanding | HEAR Sound Event Detection | DCASE2016 Task2 (Dohi et al., 2022) |
| Audio / Music Classification / Emotion | Emotion Classification in Songs | EMOTIFY (Aljanaki et al., 2016) |
| Audio / Music Classification / Emotion | MARBLE Emotion Detection | MTG MoodTheme (Bogdanov et al., 2019) |
| Audio / Rhythm Analysis | Music Beat Tracking | ASAP (Foscarin et al., 2020) |
| Audio / Rhythm Analysis | MARBLE Beat Tracking | GTZAN-Rhythm (Tzanetakis & Cook, 2002) |
| Audio / Harmony & Pitch / Pitch | Pitch Extraction by Lyrics | Children’s Song Dataset (CSD) (Choi et al., 2020) |
| Audio / Harmony & Pitch / Pitch | HEAR Music Transcription | MAESTRO Hawthorne et al. (2019) |
| Audio / Harmony & Pitch / Pitch | HEAR Percussion Instruments Tonic Classification | Mridangam Tonic (Anantapadmanabhan et al., 2013) |
| Audio / Sound Event / Environment | Emergency Traffic Detection | ETD (tabarka rajab, 2021) |
@@
| Audio / Sound Event / Environment | Environmental Sound Classification | ESC50 Natural Soundscapes And Water Sounds (Piczak, 2015) |
@@
| Audio / Music Classification / Instrument | Instrument Classification | Beijing Opera Instrument (Tian et al., 2014) |
| Audio / Music Classification / Instrument | HEAR Percussion Instruments Classification | Beijing Opera Percussion (Tian et al., 2014) |
| Audio / Music Classification / Instrument | MARBLE Instrument Classification | MTG Instrument (Bogdanov et al., 2019) |
| Audio / Sound Event / Animal | Animal Classification | Wave Source Test (WavSource, n.d.) |
| Audio / Sound Event / Animal | Cat Emotion Classification | Cat Sound Classification Dataset V2 (Pandeya et al., 2018; Pandeya & Lee, 2018) |
| Audio / Sound Event / Animal | Environmental Sound Classification | ESC50 Animals (Piczak, 2015) |
| Audio / Music Classification / Genre | HEAR Music Genre Classification | GTZAN Genre (Tzanetakis & Cook, 2002) |
| Audio / Music Classification / Genre | MARBLE Music Tagging | MTG Top50 (Bogdanov et al., 2019), MagnaTagATune (Law et al., 2009) |
| Audio / Music Classification / Genre | MARBLE Genre Classification | MTG-Genre (Bogdanov et al., 2019), GTZAN (Tzanetakis & Cook, 2002) |
| Audio / Signal-Characteristics Analysis | Sound Effect Detection | IDMT-SMT-Audio-Effects (Stein, 2010) |
| Audio / Signal-Characteristics Analysis | Audio Duration Prediction | NTUML2021 (Yang et al., 2024a) |
| Audio / Signal-Characteristics Analysis | HEAR Music Speech Classification | GTZAN Music Speech (Tzanetakis, 1999) |
| Audio / Signal-Characteristics Analysis | Speech Detection | LJSpeech (Ito & Johnson, 2017) |
| Audio / Signal-Characteristics Analysis | Speech Detection | LibriSpeech (Panayotov et al., 2015) |

### v2

@@ 4.1 Models
We evaluated several publicly-available models on Dynamic-SUPERB Phase-2: SALMONN, LTU-AS, Qwen-Audio-Chat, Qwen2-Audio-7B-Instruct, WavLLM, MU-LLaMA (Liu et al., 2024), GAMA (Ghosh et al., 2024). SALMONN is further categorized into 7B and 13B versions based on the size of its LLM component. All of these models are publicly available, and we utilized their official implementations for inference without any modifications. The first five models are instruction-based speech models, each also trained with audio understanding capabilities. However, as they were not trained on music data, we do not expect them to perform well on music-related tasks. Hence, we further included MU-LLaMA and GAMA. MU-LLaMA is specifically designed for various music-understanding tasks, and GAMA is trained with both general audio data and a music corpus. Similar to the speech models, MU-LLaMA and GAMA adopt an LLM-based framework. They use a music or audio encoder, such as MERT (LI et al., 2024) or Q-former, to convert music or audio into features, which the LLM then uses as prompts for subsequent reasoning. The sampling strategies used for generating outputs from the LLMs were retained as defined in their official implementations. Finally, we implemented a cascaded system baseline called Whisper-LLaMA. The system first transcribes audio using Whisper-v3-large, and then LLaMA3.1-8B processes the transcriptions to perform various tasks based on the provided instructions.
When evaluating models on these diverse tasks, several challenges inevitably arise. One challenge was testing the baseline models on tasks involving multiple audio inputs. For example, in speaker verification, a model determines whether two utterances are produced by the same speaker. Among the evaluated models, only Qwen-Audio and Qwen2-Audio provide interfaces that allow inputting multiple audio files. For the remaining models, we concatenated all audio files, separated by 0.5 seconds of silence, and used the concatenated file as input.
Last, different models have their own maximum supported audio durations. Models using Whisper to encode speech face an inherent limitation: Whisper can process audio files of up to 30 seconds, and some models are restricted to handling even shorter durations. Consequently, we retained the original model settings and did not modify the model architecture to accommodate longer audio clips. A preliminary analysis of all audio (Appendix G) in Dynamic-SUPERB Phase-2 shows that only a small proportion (around 6.57%) exceeds 30 seconds in length. Hence, we believe our settings do not largely impact the evaluation results and ensure reasonable inference efficiency.
@@ 4.2 Evaluation metrics
The outputs of universal speech models are natural language sentences, making it difficult to assess their correctness using conventional metrics. For classification tasks, such as emotion recognition, a task-specific model outputs a label from predefined emotions in the dataset, while a universal model generates sentences like “The speaker sounds happy.” In this case, we can easily assess the correctness of the former by comparing labels, but this approach does not apply to the latter. Similarly, in regression tasks, a natural language response like “Using MOS scoring criteria, I give the audio a score of 3 out of 5” makes it difficult to directly use metrics such as mean square error.
@@
In evaluation, we categorize tasks into three types: (1) classification, (2) regression, and (3) sequence generation. For each type, we use different pipelines to evaluate the models’ outputs. For classification tasks, we utilize external LLMs (GPT-4o) as referees, with the temperature set to be zero for evaluation consistency, to evaluate whether the outputs from speech or music models222To avoid ambiguity, unless specified otherwise, the terms “output” or “model output” in this section refer to the results generated by the models in Sec. 4.1. match the ground truth. This approach has been widely adopted in the NLP community (Wang et al., 2023; Liu et al., 2023b; Chiang & Lee, 2023), and we extend its application to speech research. We design a prompt that includes the task instructions, the model’s output to be evaluated, and the corresponding ground truth. The LLM judge gets this prompt and determines if the output aligns with the ground truth using a chain-of-thought reasoning process. By leveraging the strong instruction-following capabilities of these LLMs, we constrain them to provide the final decision in a fixed format, allowing us to extract the answer using simple methods such as regular expressions333We have tried to ask the speech universal models to directly output their decisions in a fixed format, but they do not always follow the instructions. Therefore, the evaluation pipeline described here is necessary.. We then define accuracy as the percentage of outputs that the LLM judge considers aligned with the ground truth. Please refer to Appendix E for details on the prompts and the alignment between LLMs and humans. Importantly, although we used GPT-4o for evaluation in the main text, to support reproducibility we have also included comprehensive evaluation results using the open-source LLM (LLaMA3.1-8b-Instruct) in Appendix C.
@@
For regression tasks, the above method cannot be applied because performance is not assessed using hard labels. Instead, we use LLMs (GPT-4o) as a post-processor to transform natural language responses into a format compatible with the original metrics used in these tasks. Specifically, we developed a prompt that directs the LLMs to extract essential information from the output and convert it into the same format as the ground truth. If the output lacks correct or relevant information that can be converted to match the ground truth format, the LLM returns “N/A” to mark it as invalid. Since conventional metrics cannot accommodate “N/A” and setting a default value is unreasonable due to the variability of metrics, we also calculate the N/A rate for each regression task, defined as the percentage of invalid outputs within a task. A higher N/A rate indicates that the model struggles with following instructions, which may indirectly affect its performance on the task.
@@ 5.1 Domain-Level Performance Comparison in Taxonomy
Figure 3 presents a domain-level comparison of relative scores across different models. Whisper-LLaMA consistently has zero scores throughout all domains, serving as the reference baseline for evaluating other models. In speech domains, no model outperforms Whisper-LLaMA in speech recognition and spoken language understanding. This shows that using ASR remains a strong baseline for language understanding, as text more explicitly represents semantic information than speech. However, Whisper-LLaMA lags behind SALMONN in phoneme recognition, a task for which SALMONN is explicitly trained, indicating that task-specific training still provides superior performance in specialized domains.
Conversely, in the speaker and paralinguistics domains, all models surpass the baseline. This improvement occurs because the ASR process in the cascaded system tends to discard critical information from the speech signal, such as speaker characteristics, pitch, and emotion. In contrast, other models utilize soft representations that retain more speech information, giving them the potential to learn beyond ASR, and some of these models were also trained to perform specific tasks within certain domains 444Although the models have been trained on some tasks in the same domains, the training tasks of these models generally do not use the same instructions as Dynamic-SUPERB and do not offer as broad coverage.. Additionally, music-specific large language models like GAMA-IT and Mu-LLaMA perform poorly on speech recognition and understanding tasks, which is expected since they were primarily designed for music understanding. Nevertheless, they surprisingly achieved scores comparable to the speech models in speaker and paralinguistics.
Turning to audio and music domains, where ASR models cannot transcribe non-speech information into text, most models outperform the baseline in several areas, such as music classification and sound event detection. Surprisingly, spoken language models primarily trained on speech data (Qwen models, SALMONN, WavLLM)555Qwen and SALMONN are trained on a mixture of speech (dominant component), music, and audio data. outperform the two music models (GAMA-IT and Mu-LLaMA) in various music-related domains such as harmony and pitch, music classification, and rhythm analysis. We speculate that training on diverse data, even with significant differences in signal-level characteristics, enhances performance across domains, emphasizing the importance of developing unified models for speech, music, and general audio processing.
@@ 5.2 Core tasks results
| Whisper-LLaMA | 100.1 | 36.5 | 36.5 | 12.5 | 34.0 | 46.5 | 53.6 | 51.3 | 45.0 | 1068.7 |
| SALMONN-7B | 25.4 | 30.5 | 21.0 | 12.1 | 15.1 | 49.0 | 61.6 | 61.5 | 45.0 | - |
| SALMONN-13B | 24.6 | 2.0 | 5.5 | 12.5 | 2.8 | 51.5 | 29.6 | 106.3 | 51.0 | - |
| Qwen-Audio-Chat | 100.6 | 60.5 | 26.5 | 70.4 | 69.4 | 48.0 | 42.2 | 90.1 | 49.0 | 8852.2 |
| Qwen2-Audio-7B-Inst. | 101.0 | 47.0 | 26.0 | 75.8 | 36.7 | 53.5 | 42.4 | 55.7 | 51.0 | 4664.4 |
| WavLLM | 100.0 | 43.0 | 28.0 | 12.1 | 6.9 | 49.5 | 50.8 | 84.9 | 49.0 | 3621.2 |
| LTU-AS | 102.8 | 1.0 | 0.1 | 25.8 | 96.3 | 45.0 | 10.2 | 559.3 | 44.0 | 14923.3 |
| GAMA-IT | 100.4 | 2.0 | 1.0 | 0.0 | 116.7 | 2.5 | 2.7 | 4750.0 | 45.0 | 187.6 |
| Mu-LLaMA | 110.3 | 4.0 | 3.5 | 12.9 | 103.7 | 51.0 | 1.0 | 15869.6 | 44.0 | - |
@@
We tested the performance of the models on core tasks that are widely studied and commonly used. Table 2 presents the detailed results of the models on SUPERB tasks, displaying each task’s performance in its own metric rather than relative scores. Please refer to Appendix D for results on HEAR and MARBLE. It is worth noting that these tasks have been adapted to the Dynamic-SUPERB framework. Therefore, the results are not directly comparable to previous studies that tested on the original SUPERB. However, they indicate current performance levels on these tasks.
No single model excels across all tasks. In phoneme recognition (PR), the SALMONN models were the only ones to achieve a relatively lower phoneme error rate (PER) compared to the others. For keyword spotting (KS), Qwen-Audio-Chat was the only model to perform slightly better than a random guess (50%). Whisper-LLaMA surpassed all other models in intent classification (IC). Regarding emotion recognition (ER), Qwen-Audio-Chat and Qwen2-Audio-7B-Instruct stood out, with the latter even achieving about 75% accuracy. In ASR, SALMONN-13B and WavLLM are the only two models that achieved a word error rate (WER) lower than 10%. Query-by-example (QbE) appears to be very challenging for all models, as none performed better than a random guess (50%). Slot-filling (SF) is evaluated using two metrics: F1 for slot type and CER for slot value. While Whisper-LLaMA and SALMONN-7B are the top two in this task, their results are still far from ideal. In speaker verification (SV), all models performed badly, as even the best ones achieved results close to random guessing (50%). In speaker diarization (SD), results were similarly poor, with some even reaching a 100% N/A rate. Comparing Figure 3 and Table 2, we observed that performance on core tasks sometimes deviates from trends observed at the domain level. For example, although SALMONN-13B outperforms other models in ASR, it lags behind Whisper-LLaMA in the speech recognition domain, revealing its limited capabilities and highlighting the need to enhance its generalizability. In summary, among tested models, some excel in specific tasks, while others perform poorly across the board, and none dominate across all tasks. We believe these findings provide valuable insights for researchers aiming to improve models, starting from core tasks and progressively tackling each domain in the benchmark for broader applicability.
@@ 6 Conclusions
This paper presents Dynamic-SUPERB Phase-2, the largest benchmark for evaluating instruction-based universal spoken language models, building upon collaborative efforts across the research community. Dynamic-SUPERB covers a wide range of diverse tasks and offers a fine-grained task taxonomy. The recent models show good performance on specific tasks but poor generalization across common tasks like those in SUPERB, highlighting the need for further research on universal models. All materials are made openly available to facilitate reproduction and benchmarking. We sincerely invite researchers to join our vibrant community and collaborate to advance the field.
@@ 7 Ethics Statement
Dynamic-SUPERB Phase-2 includes several datasets, and we asked contributors to describe how they used them. Most contributors derived task data by uniformly sampling from the original datasets without applying any special processing. Therefore, Dynamic-SUPERB Phase-2 inevitably inherits the biases present in these datasets across tasks. Additionally, we asked all contributors to provide the necessary license information to ensure that we can include them in the benchmark properly. More specifically, all task data in Dynamic-SUPERB are derived from datasets with licenses that permit remixing and redistribution. As we expect the benchmark to grow dynamically in the future, we maintain a complete list of dataset licenses on our official GitHub page (https://github.com/dynamic-superb/dynamic-superb/blob/main/docs/dataset_license.md), which will be updated as new tasks are introduced. If there are any changes to the dataset license in the future, we may adjust the benchmark task accordingly.
@@ 8 Reproducibility Statement
We open-source all task data and the evaluation pipeline at https://github.com/dynamic-superb/dynamic-superb. For LLM-based evaluation, we set GPT-4o’s temperature to 0 to ensure a stable evaluation process. However, we acknowledge that GPT-4o’s behavior may change with future updates. Therefore, we provide a comprehensive set of evaluation results using LLaMA-3.1-8b-Instruct in Appendix C (Table LABEL:tab:eval-result-llama). Additionally, in Appendix E, we present a preliminary study on using open-source LLMs for evaluation, demonstrating promising correlations with human annotations. We encourage researchers to use these open-source LLMs when evaluating their own models to support reproducibility.
9 Acknowledgement
We thank the National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.
@@ Appendix B Dynamic-SUPERB tasks
| Domain | Task | Dataset |
| Audio / Harmony & Pitch / Harmony | Chord Classification | Musical Instrument Chord Classification (Audio) (DeepContractor, 2024) |
| Audio / Harmony & Pitch / Harmony | MARBLE Key Detection | GiantSteps (key) (Knees et al., 2015) |
| Audio / Harmony & Pitch / Pitch | HEAR Music Transcription | MAESTRO (Hawthorne et al., 2019) |
| Audio / Harmony & Pitch / Pitch | HEAR Percussion Instruments Tonic Classification | Mridangam Stroke (Anantapadmanabhan et al., 2013) |
| Audio / Harmony & Pitch / Pitch | Pitch Extraction By Lyrics | CSD (Choi et al., 2020) |
| Audio / Music Classification / Emotion | Emotion Classificaiton In Songs | EMOTIFY (Aljanaki et al., 2016) |
| Audio / Music Classification / Emotion | MARBLE Emotion Detection | MTG-Jamendo (Bogdanov et al., 2019) |
| Audio / Music Classification / Genre | HEAR Music Genre Classification | ISMIR04 (Cano et al., 2018) |
| Audio / Music Classification / Genre | MARBLE Genre Classification MTG-Genre | MTG-Jamendo (Bogdanov et al., 2019) |
| Audio / Music Classification / Genre | MARBLE Music Tagging | MagnaTagATune (Law et al., 2009) |
| Audio / Music Classification / Genre | MARBLE Music Tagging | MTG-Jamendo (Bogdanov et al., 2019) |
| Audio / Music Classification / Instrument | HEAR Percussion Instruments Classification | Beijing Opera Percussion Instrument Dataset (Tian et al., 2014) |
| Audio / Music Classification / Instrument | MARBLE Instrument Classification | MTG-Jamendo (Bogdanov et al., 2019) |
| Audio / Rhythm Analysis | MARBLE Beat Tracking ASAP | ASAP (Foscarin et al., 2020) |
| Audio / Rhythm Analysis | Music Beat Tracking ASAP | ASAP (Foscarin et al., 2020) |
| Audio / Safety / Audio Integrity | Audio Editing Identification | People’s Speech (Galvez et al., 2021) |
| Audio / Safety / Audio Integrity | Scene Fake Detection | ASPIRE (Gogate et al., 2020) |
| Audio / Safety / Deepfake | Audio Deep Fake Detection | LJSpeech (Ito & Johnson, 2017) WaveFake (Frank & Schönherr, ) MUSDB18HQ (Rafii et al., 2019) |
| Audio / Safety / Deepfake | Singing Voice Deepfake Detection | CtrSVDD (Zang et al., 2024) ACEKiSing (Shi et al., 2024a) M4Singer (Zhang et al., 2022) |
| Audio / Signal Characteristics Analysis | Audio Duration Prediction | NTUML2021 (Yang et al., 2024a) |
| Audio / Signal Characteristics Analysis | HEAR Music Speech Classification | MAESTRO (Hawthorne et al., 2019) Librispeech (Panayotov et al., 2015) |
| Audio / Signal Characteristics Analysis | Sound Effect Detection | RemFx (Rice et al., 2023) |
| Audio / Signal Characteristics Analysis | Speech Detection | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Audio / Signal Characteristics Analysis | Speech Detection | LibriSpeech-TestOther (Panayotov et al., 2015) |
| Audio / Signal Characteristics Analysis | Speech Detection | LJSpeech (Ito & Johnson, 2017) |
| Audio / Singing Analysis / Lyrics | Lyric Translation | SingSet (self-recorded) |
| Audio / Singing Analysis / Lyrics | Song Lyric Recognition | SingSet (self-recorded) |
@@
| Audio / Sound Event / Advanced Understanding | HEAR Sound Event Detection | DCASE2016Task2 (IEEE DCASE 2016 Challenge, 2016) |
| Audio / Sound Event / Advanced Understanding | Multi-channel Sound Event Understanding | STARSS23 (Shimada et al., 2024) |
| Audio / Sound Event / Animal | Animal Classification | WaveSource-Test (WavSource, n.d.) |
| Audio / Sound Event / Animal | Cat Emotion Classification | CatSoundClassificationDataset-V2 (Pandeya et al., 2018; Pandeya & Lee, 2018) |
| Audio / Sound Event / Animal | Environmental Sound Classification | ESC50-Animals (Piczak, 2015) |
@@
| Audio / Sound Event / Environment | Emergency Traffic Detection | Large-Scale-Audio-dataset (Asif et al., 2022) |
@@
| Audio / Sound Event / Environment | Environmental Sound Classification | ESC50-NaturalSoundscapesAndWaterSounds (Piczak, 2015) |
| Audio / Spatial Audio | Audio Spatial Distance Prediction | Spatial LibriSpeech (Sarabia et al., 2023) |
| Audio / Spatial Audio | How Far Are You | 3DSpeaker (Zheng et al., 2023) |
| Audio / Spatial Audio | Sound Position Prediction | BinauralSoundPerception (Vasudevan et al., 2020) |
| Speech / Paralinguistics / Emotion Analysis | Emotion Recognition | MELD (Poria et al., 2019) |
| Speech / Paralinguistics / Emotion Analysis | SUPERB Emotion Recognition RAVDESS | RAVDESS (Livingstone & Russo, 2018) |
| Speech / Paralinguistics / Vocal Event Detection | Human Non-Speech Sound Recognition | Nonspeech7k-test (Rashid et al., 2023) CommonVoice (Ardila et al., 2020) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Phoneme Recognition Tasks | SUPERB Phoneme Recognition | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Phoneme Recognition Tasks | Phoneme Recognition | LibriSpeech-TestOther (Panayotov et al., 2015) |